OpenStudy (mindblast3r):

78 is 15% of what number?

OpenStudy (paki):

78=15/100*X... solve for X now...

OpenStudy (mindblast3r):

how do i solve for x??

OpenStudy (mindblast3r):

turn 15% into a decimal?

OpenStudy (mindblast3r):

i actually understand how to get the answer it's 520, but I'm not really sure how it works.

OpenStudy (paki):

simple... X=78*100/15... now solve it....

OpenStudy (mindblast3r):

why did you multiply it by 100?

OpenStudy (paki):

please be sober.... and try not to use abusive language....

OpenStudy (mindblast3r):

sorry.

OpenStudy (paki):

100 means percent.... we are dealing here with percent... so that's why... got it now...

OpenStudy (mindblast3r):

so when we multiply 78 by 100 what does it do?

OpenStudy (mindblast3r):

7800

OpenStudy (paki):

X=7800/15

OpenStudy (mindblast3r):

7800 = 100%?

OpenStudy (paki):

yeah,,,, now divide it by 15,,, that will be the answer....

OpenStudy (mindblast3r):

are you sure? when i do .15 x 7800 = 1170

OpenStudy (paki):

X=78*100/15 x=7800/15 x=520